OpenStudy (anonymous):

Please help give me an answer? A country's population in 1993 was 72 million. In 1996 it was 76 million. Estimate the population in 2003 using exponential growth.

OpenStudy (anonymous):

easy way or hard way?

OpenStudy (anonymous):

Easy :P

OpenStudy (anonymous):

lets do the easy way first

OpenStudy (anonymous):

put 1993 as year 0 1996 as year 3 and 2003 as year 10

OpenStudy (anonymous):

the just calculate \[72\times (\frac{76}{72})^{\frac{10}{3}}\] and you will have your answer

OpenStudy (anonymous):

we can redo this the hard way if you like

OpenStudy (anonymous):

Hang on

OpenStudy (anonymous):

ok i will check the answer

OpenStudy (anonymous):

i can also explain why this simple method works if you like

OpenStudy (anonymous):

I'm not getting an answer from that equation...

OpenStudy (anonymous):

Sure, I just really need help so I can have a quick answer...

OpenStudy (anonymous):

86

OpenStudy (anonymous):

just type it in a calculator

OpenStudy (anonymous):

i did, I have a messed up calculator though and it gave me errors.. Would you mind helping me on another too..?

OpenStudy (anonymous):

you have to be careful with parentheses

OpenStudy (anonymous):

sure ask away

OpenStudy (anonymous):

A country's population in 1994 was 195 million. In 2002 it was 199 million. Estimate the population in 2016 using exponential growth.

OpenStudy (anonymous):

ok we go right to the answer

OpenStudy (anonymous):

1994 is 0 2002 is 8 2016 is 22

OpenStudy (anonymous):

we compute \[195\times (\frac{199}{195})^{\frac{22}{8}}\]

OpenStudy (anonymous):

The exponent is 22 over what...? I can't read it.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

answer is 206

OpenStudy (anonymous):

is it clear where i am getting these numbers from?

OpenStudy (anonymous):

Kind of, yes, I just wish my calculator wasn't messed up.

OpenStudy (anonymous):

you have to be careful with the calculator but since you are already on line just use this http://www.wolframalpha.com/input/?i=195%28199%2F195%29^%2822%2F8%29

OpenStudy (anonymous):

I have one last one if you could help me? it's easier than those, but I would just like to make sure. :)

OpenStudy (anonymous):

knock yourself out

OpenStudy (anonymous):

A culture strand started with 6000. After 4 hours, it grew to 6,600 bacteria. Predict how many bacteria will be present after 10 hours.

OpenStudy (anonymous):

yes this is easier. we don't need to adjust for the year.

OpenStudy (anonymous):

also \[\frac{6600}{6000}=1.1\] so put \[6000\times(1.1)^{\frac{10}{4}}\]

OpenStudy (anonymous):

ready for an explanation or is the answer enough?

OpenStudy (anonymous):

Hang on I will try that equation.

OpenStudy (anonymous):

type it in to wolframalfpha just as i did the last one. very user friendly

OpenStudy (anonymous):

7614?

OpenStudy (anonymous):

It was a decimal but I rounded it to a whole number. Is that correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:D Awesome, thank you so much!

OpenStudy (anonymous):

this is not how your math teacher wants you to do it, believe me. but it is the easiest way

OpenStudy (anonymous):

Well.. I do not have any math teachers. long story.

OpenStudy (anonymous):

let me know if you want either an explanation or the math teacher method

OpenStudy (anonymous):

on line? home school?

OpenStudy (anonymous):

I have to take and finish a whole years worth of alg 2 and some pre-calc in 3 weeks. I never took any of these classes, but I can't get my scholarship sent to my college without a completion grade. :(

OpenStudy (anonymous):

keep coming here then

OpenStudy (anonymous):

Thanks :)